### v1

# DMS*: Minimizing Makespan for Multi-Agent Combinatorial Path Finding

Abstract
Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their initial to goal locations, while visiting a set of intermediate target locations in the middle of the paths. MCPF is challenging as it involves both planning collision-free paths for multiple agents and target sequencing, i.e., solving traveling salesman problems to assign targets to and find the visiting order for the agents. Recent work develops methods to address MCPF while minimizing the sum of individual arrival times at goals. Such a problem formulation may result in paths with different arrival times and lead to a long makespan, the maximum arrival time, among the agents. This paper proposes a min-max variant of MCPF, denoted as MCPF-max, that minimizes the makespan of the agents. While the existing methods (such as MS*) for MCPF can be adapted to solve MCPF-max, we further develop two new techniques based on MS* to defer the expensive target sequencing during planning to expedite the overall computation. We analyze the properties of the resulting algorithm Deferred MS* (DMS*), and test DMS* with up to 20 agents and 80 targets. We demonstrate the use of DMS* on differential-drive robots.
Index Terms:
Path Planning for Multiple Mobile Robots or Agents, Multi-Agent Path Finding, Traveling Salesman Problem
I Introduction
Multi-Agent Path Finding (MAPF) seeks a set of collision-free paths for multiple agents from their respective start to goal locations. This paper considers a generalization of MAPF called Multi-Agent Combinatorial Path Finding (MCPF), where the agents need to visit a pre-specified set of intermediate target locations before reaching their goals. MAPF and MCPF arise in applications such as surveillance [1] and logistics [2]. For instance, factories [3, 4] use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing. These robots share a cluttered environment and follow collision-free paths. In such settings, MAPF problems and their generalizations naturally arise to optimize operations.
MCPF is challenging as it involves both collision avoidance among the agents as present in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [5, 6] to specify the allocation and visiting orders of targets for all agents. Both the TSP and the MAPF are NP-hard to solve to optimality [7, 5], and so is MCPF. A few methods have been developed [8, 9] to address the challenges in MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times. Such a formulation may result in an ensemble of paths where some agents arrive early while others arrive late, which leads to long execution times before all agents finish their paths. This paper thus proposes a min-max variant of MCPF (Fig. 1), denoted as MCPF-max, where the objective is to minimize the maximum arrival time, which is also called the makespan, of all agents.
To solve MCPF-max, this paper first adapts our prior MS* algorithm [8], which was first designed for MCPF-sum, to address MCPF-max. Then, we further develop two new techniques to expedite the planning, and we call the resulting new algorithm Deferred MS* (DMS*).
Figure 1: MCPF-max problem, which seeks a set of collision-free paths while minimizing the maximum arrival time of all agents. The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal. The orange target appears on the path of the blue agent, but is not claimed to be visited by the blue agent.
Specifically, the existing MS* follows the general heuristic search framework (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state. MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for agents. When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs for the agents to arrive at their goals. Therefore, the cost of target sequences provides an admissible heuristic to guide the state selection and expansion as in A*. Furthermore, MS* leverages the idea in M* [10] to first use the target sequences to build a low-dimensional search space, and then grow this search space by coupling agents together for planning only when collision happens. By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques under the heuristic search framework and is able to provide completeness and solution quality guarantees.
The first technique developed in this paper is applicable to MS* for both MCPF-max and MCPF-sum. When expanding a state, a set of successor states are generated, and for each of them, MS* needs to invoke the TSP solver to find the target sequence and the heuristic value of this successor state. Since the number of successor states can be large for each expansion, MS* needs to frequently invokes the TSP solver which slows down the computation. To remedy this issue, for each generated successor, we first use a fast-to-compute yet roughly estimated cost-to-go as the heuristic value, and defer calling TSP solver for target sequencing until that successor is selected for expansion.
The second technique developed in this paper is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum. Since the goal here is to minimize the makespan, during the search, agents with non-maximum arrival time naturally have “margins” in a sense that they can take a longer path without worsening the makespan of all agents. We take advantage of these margins to let agents re-use their previously computed target sequences and defer the expensive calls of TSP solvers until the margin depletes.
To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as real robot experiments. The simulation shows that the new techniques in DMS* help triple the success rates and reduce the average runtime to solution. The real robot experiments demonstrate that the planned path are executable on real robots, and that DMS* is able to optimize makespan while the existing planners that minimize the sum of arrival times can lead to long makespan.
II Related Work
II-A Multi-Agent Path Finding
To find collision-free paths for multiple agents, a variety of MAPF algorithms were developed, falling on a spectrum from coupled [11] to decoupled [12], trading off completeness and optimality for scalability. In the middle of this spectrum lie the dynamically-coupled methods such as M* [10] and CBS [13], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary in order to resolve agent-agent collision. These dynamically-coupled methods were improved and extended in many ways [14, 15, 16, 17].
II-B Traveling Salesman Problems
In the presence of multiple intermediate targets, the planner needs to determine both the assignment and visiting order of the targets for the agents. For a single agent, the well-known Traveling Salesman Problem (TSP), which seeks a shortest tour that visits every vertex in a graph, is one of the most well-known NP-hard problems [5]. Closely related to TSP, the Hamiltonian Path Problem (HPP) requires finding a shortest path that visits each vertex in the graph from a start vertex to a goal vertex. The multi-agent version of the TSP and HPP (denoted as mTSP and mHPP, respectively) are more challenging since the vertices in the graph must be allocated to each agent in addition to finding the optimal visiting order of vertices. To simplify the presentation, we refer to all these problems simply as TSPs. To solve TSPs, a variety of methods have been developed ranging from exact methods (branch and bound, branch and price) [5] to heuristics [18] and approximation algorithms [19], trading off solution optimality for runtime efficiency. This work does not develop new TSPs solvers and leverage the existing ones.
II-C Target Assignment, Sequencing and Path Finding
Some recent research combines MAPF with target assignment and ordering [20, 21, 22, 23, 24, 25, 9, 8, 26, 27, 28, 29]. Most of them either consider target assignment only (without the need for computing visiting orders of targets) [20, 21, 22], or consider the visiting order only given that each agent is pre-allocated a set of targets [23, 24, 25]. Our prior work [8, 9] seeks to handle the challenge in target assignment and ordering as well as the collision avoidance in MAPF simultaneously. These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max, which seeks to minimize the maximum arrival time.
III Problem
Let index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents. All agents share a workspace that is represented as an undirected graph $G^{W}=(V^{W},E^{W},c^{W})$, where $W$ stands for workspace. Each vertex $v\in V^{W}$ represents a possible location of an agent. Each edge $e=(u,v)\in E^{W}\subseteq V^{W}\times V^{W}$ represents an action that moves an agent between $u$ and $v$. $c^{W}:E^{W}\rightarrow(0,\infty)$ maps an edge to its positive cost value. In this paper, the cost of an edge is equal to its traversal time, and each edge has a unit cost. [^1] In [8, 9], the edge cost can be different from its traversal time: the edge cost can be any positive number while the traversal time is always one unit per edge. In this paper, the edge cost are the same as the edge traversal time, which is one for each edge.
Let the superscript $i\in I$ over a variable denote the specific agent to which the variable belongs (e.g. $v^{i}\in V^{W}$ means a vertex corresponding to agent $i$). Let $v_{o}^{i},v_{d}^{i}\in V^{W}$ denote the initial (or original) vertex and the goal (or destination) vertex of agent $i$ respectively. Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}$ denote the set of target vertices. For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.
Let $\pi^{i}(v^{i}_{1},v^{i}_{\ell})$ denote a path for agent $i$ between vertices $v^{i}_{1}$ and $v^{i}_{\ell}$, which is a list of vertices $(v^{i}_{1},v^{i}_{{2}},\dots,v^{i}_{\ell})$ in $G^{W}$ with $(v^{i}_{k},v^{i}_{k+1})\in E^{W},k=1,2,\cdots,\ell-1$. Let $g(\pi^{i}(v^{i}_{1},v^{i}_{\ell}))$ denote the cost of the path, which is the sum of the costs of all edges present in the path: $g(\pi^{i}(v^{i}_{1},v^{i}_{\ell}))=\Sigma_{j=1,2,\dots,{\ell-1}}c^{W}(v^{i}_{{j}},v^{i}_{{j+1}})$.
All agents share a global clock and start to move along their paths from time $t=0$. Each action of the agents, either wait or move along an edge, requires one unit of time. Any two agents $i,j\in I$ are in conflict if one of the following two cases happens. The first case is a vertex conflict $(i,j,v,t)$ where two agents $i,j\in I$ occupy the same vertex $v$ at the same time $t$. The second case is an edge conflict $(i,j,e,t)$, where two agents $i,j\in I$ go through the same edge $e$ from opposite directions between times $t$ and $t+1$.
Definition 1 (MCPF-max Problem).
The Multi-Agent Combinatorial Path Finding with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited [^2] In MCPF-max, the notion that an agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited. In other words, if a target $v$ is in the middle of the path of agent $i$ and agent $i$ does not claim $v$ is visited, then $v$ is not considered as visited. Additionally, a visited target $v$ can appear in the path of another agent. For the rest of the paper, when we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target. at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}cost(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
Remark 1.
The MCPF-max is similar to MCPF-sum [9, 8]. The only difference between these two problems is that: MCPF-sum seeks to minimize the sum of individual path costs while MCPF-max minimizes the maximum of individual path costs.
IV Method
This section begins with a toy example in Fig. 2 to illustrate the planning process of DMS*, and provides the explanation in the caption. This section then introduces the concepts and notations, explains the pseudo-code and elaborates the technical detail.
Figure 2: An illustration of DMS* and related concepts. (a) shows the workspace graph $G^{W}$ and a toy problem. (b) shows the corresponding target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices. (c) shows the workflow of DMS* in Alg. 1. (d) shows that DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1). This joint sequence $\gamma(l_{o})$ can be converted to a joint path $\pi(l_{o})$ (d2,d3), whose corresponding makespan $f_{max}$ is 9. (e) The joint path $\pi(l_{o})$ leads to a policy $\phi_{\gamma(l_{o})}$ that maps a joint vertex to the next joint vertex along $\pi(l_{o})$. Since agent-agent conflicts are ignored, when following the policy, agents may run into conflict (e2). When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the nearby ancestor labels. (f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of the agents that are in conflict (f1,f2,f3). During this re-expansion, for each of the successor, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP. For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means that each agent can move at most one step towards their goals. When either of (f2) and (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and see if the popped label should be expanded or re-added to OPEN for future expansion. To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this will worsen the overall makespan. For the successor shown in (f2), the resulting makespan is 9 (f4), which is no worse than the previous makespan 9, and DMS* will not call mHPP solver to save computational effort. For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find an optimal joint sequence from that successor (f3) to the goals. Finally, with the joint sequence, new policy for the successors can be built and the search continues as in (e).
IV-A Concepts and Notations
IV-A1 Joint Graph
Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of all agents which is the Cartesian product of $N$ copies of $G^{W}$, where each vertex $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices. The joint vertex $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ corresponding to the initial vertices of all agents is called the initial joint vertex. Let $\pi(u,v),u,v\in\mathcal{V^{W}}$ denote a joint path, which is a tuple of $N$ (individual) paths of the same length, i.e., $\pi(u,v)=(\pi^{1}(u^{1},v^{1}),\cdots,\pi^{N}(u^{N},v^{N}))$. For the rest of this paper, we use a “path” to denote an “individual path”. DMS* searches the joint graph $\mathcal{G^{W}}$ for a joint path that solves the MCPF-max.
IV-A2 Binary Vector
For any joint vertex $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g. $\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited along the joint path, and we need to differentiate between them during the search. We therefore introduce the following concepts. First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{t_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list. [^3] In this paper, for a vector related to targets (e.g. a binary vector $\vec{a}$ of length $M$), we use subscripts (e.g. $a_{m}$) to indicate the elements in the vector. For a vector or joint vertex that is related to agents (e.g. $\vec{g}$ of length $N$, or $v\in\mathcal{V^{W}}$), we use a superscript (e.g. $i$ in $g^{i},v^{i}$) to indicate the element in the vector or joint vertex corresponding to an agent. Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if the corresponding target $t_{m}$ is visited by some agent, and $a_{m}=0$ otherwise.
IV-A3 Label
Let $l=(v,\vec{a},\vec{g})$ denote a label, where $v$ is a joint vertex, $\vec{a}$ is a binary vector of length $M$, and $\vec{g}$ is a cost vector of length $N$. Here, each component $g^{i},i\in I$ corresponds to the path cost of agent $i$. Intuitively, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path costs as specified by $\vec{g}$. To simplify the presentation, given a label $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$. Furthermore, let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$. To solve the MCPF-max in Def. 1, the algorithm needs to find a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reach the goals, and $g_{max}(l)$ is the objective value to be minimized.
IV-A4 Label Comparison
To compare two labels at the same joint vertex (i.e., to compare the joint paths represented by these two labels), we need to compare both their binary vectors as well as the path costs. We introduce the following dominance relationship to compare two binary vectors.
Definition 2 (Binary Dominance).
For any two binary vectors $\vec{a}$ and $\vec{b}$, $\vec{a}$ dominates $\vec{b}$ ($\vec{a}\succeq_{b}\vec{b}$), if both the following conditions hold: (i) $\forall m\in\{1,2,\cdots,M\}$, ${a}_{m}\geq{b}_{m}$; (ii) $\exists m\in\{1,2,\cdots,M\}$, ${a}_{m}>{b}_{m}$.
Intuitively, $\vec{a}$ dominates $\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$. In addition, two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
To compare two labels with the same joint vertex, we define the following dominance rule between them.
Definition 3 (Label Dominance).
For any two labels ${l}_{1}$ and ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either of the following two conditions holds: (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$.
Intuitively, if $l_{1}$ dominates $l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$. If $l_{1}$ does not dominate $l_{2}$, $l_{1}$ is then non-dominated by $l_{2}$. Any two labels are non-dominated (with respect to each other) if each of them is non-dominated by the other. Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if they have the same vertex, binary vector and $g_{max}$ (i.e., $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$). Note that there is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Def. III seeks to minimize the maximum path cost over the agents. Finally, for each joint vertex $v\in\mathcal{V^{W}}$, let $\mathcal{F}(v)$ denote a set of labels that are non-dominated to each other during the search.
IV-A5 Target Sequencing
Let $\gamma^{i}=\{v^{i}_{o},t_{1},t_{2},\cdots,t_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where $v^{i}_{o}\in V_{o}$ is the initial vertex of agent $i\in I$, each $t_{j}\in V_{t},j=1,2,\cdots,k$ is a target vertex and $v^{i}_{d}\in V_{d}$ is a goal vertex. A target sequence $\gamma^{i}$ specifies both the assignment and visiting order of targets for agent $i$. Let $\gamma=\{\gamma^{1},\gamma^{2},\cdots,\gamma^{N}\}$ denote a joint sequence, which specify the assignment and visiting order of all targets for all agents.
Given a label $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited. We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, which is a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$. In other words, (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$, the vertex of agent $i$ in $v(l)$, visits a set of unvisited targets $\{t_{m}\}$, where the $m$-th component of $\vec{a}(l)$ is zero, and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ collectively visits all unvisited targets as specified in $\vec{a}(l)$.
IV-A6 Target Graph
Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in the workspace graph $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$. The cost of a target sequence $cost(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$.
Given a label $l$, to find a joint sequence $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved. Let SolveMHPP($l$) denote a procedure that solves the mHPP. SolveMHPP($l$) first formulates a target graph $G^{T}=(V^{T},E^{T},c^{T})$, which is computed based on the workspace graph $G^{W}$ as follows. The vertex set $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{t_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$. $G^{T}$ is fully connected and $E^{T}=V^{T}\times V^{T}$. The edge cost in $G^{T}$ of any pair of vertices $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the workspace graph $G^{W}$. An example $G^{T}$ is shown in Fig. 2(b). The aforementioned target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$ and a joint sequence $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as specified by $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints. The procedure SolveMHPP$(l)$ can be implemented by various existing algorithms for mHPP.
IV-A7 Heuristic and Policy
Given a joint sequence $\gamma(l)$, to simplify the notations, let $h^{i}(l):=cost(\gamma^{i}(l))$ be a heuristic value. Since agent-agent conflicts are ignored along the target sequences, the vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$. When SolveMHPP($l$) solves the mHPP to optimality, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which can be used as an admissible heuristic for the search. For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from the initial vertices to goals for all agents by further extending the joint path represented by $l$. Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to label $l$.
Given a label $l$ and its joint sequence $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ can be built out of $\gamma(l)$, which maps one label to another label as follows. First, a joint path $\pi$ can be built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l)$ for any agent $i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$. Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from one joint vertex $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$. A set of corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ can be built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by considering if $v_{k}$ visits any new targets. A set of corresponding cost vectors $\vec{g}_{k}$ can also be computed in a similar way as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$. As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l_{k}=(v_{k},\vec{a}_{k},\vec{g}_{k})$ to the next label $l_{k+1}=(v_{k+1},\vec{a}_{k+1},\vec{g}_{k+1})$ along the target sequence $\phi_{\gamma(l)}$. Given a vertex $v^{i}$ of agent $i$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$. An example of $\phi^{i}_{\gamma(l)}$ is shown in Fig. 2(d). A label $l$ is called on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$ for some $l^{\prime}$. Otherwise, $l$ is called off-policy, which means its next label is unknown yet and a mHPP needs to be solved from $l$ to find the policy $\phi_{\gamma(l)}$. We denote $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) as the process of computing the joint sequence, heuristic values and joint policy.
IV-B DMS* Algorithm
Algorithm 1 Pseudocode for DMS*
[…]
2:$parent(l_{o})\leftarrow NULL$, $f_{temp}(l_{o})\leftarrow 0$
3:add $l_{o}$ to OPEN with $f_{temp}(l_{o})$ as the priority
4:add $l_{o}$ to $\mathcal{F}(v_{o})$
5:while OPEN is not empty do
6:    $l=(v,\vec{a},\vec{g})\leftarrow$ OPEN.pop()
7:    TargetSeq($parent(l)$, $l$, $I_{C}(parent(l))$)
8:    $f_{max}(l)\leftarrow\max_{i\in I}\{\vec{g}(l)+w\cdot\vec{h}(l)$}
9:    if $f_{max}(l)>f_{temp}(l)$then
10:         $f_{temp}(l)\leftarrow f_{max}(l)$
11:         add $l$ to OPEN with $f_{temp}(l)$ as the priority
12:         continue
[…]
14:    return Reconstruct($l$)
15:    $L_{succ}\leftarrow$ GetSuccessors($l$)
16:    for all $l^{\prime}\in L_{succ}$ do
[…]
18:         BackProp($l$, $I_{C}(l^{\prime})$)
19:         if $I_{C}\neq\emptyset$ then
20:             continue
21:         if IsDominated($l^{\prime}$then
22:             DomBackProp($l,l^{\prime}$)
23:             continue
[…]
25:         $parent(l^{\prime})\leftarrow l$
26:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
27:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
28:return Failure (no solution)
Algorithm 2 Pseudocode for TargetSeq($l$, $l^{\prime}$, $I_{C}(l)$)
1:INPUT: two labels $l,l^{\prime}$ where $l$ is the parent of $l^{\prime}$.
2:if $l^{\prime}$ is on-policy then
3:    $\gamma(l^{\prime})\leftarrow\gamma(l)$
4:    return
5:$\gamma(l^{\prime})\leftarrow\gamma(l)$
6:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
7:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
[…]
9:return
Algorithm 3 Pseudocode for BackProp($l,I_{C}(l^{\prime})$)
1:if $I_{C}(l^{\prime})\nsubseteq I_{C}(l)$ then
2:    $I_{C}(l)\leftarrow I_{C}(l^{\prime})\bigcup I_{C}(l)$
3:    if $l\notin$ OPEN then
4:         add $l$ to OPEN
5:    for all $l^{\prime\prime}\in$ back_set($l$do
6:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
Algorithm 4 Pseudocode for DomBackProp($l,l^{\prime}$)
1:for all $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ do
2:    if $l^{\prime\prime}\succeq_{l}l^{\prime}\text{ or }l^{\prime\prime}=l^{\prime}$ then
3:         BackProp($l$, $I_{C}(l^{\prime\prime})$)
4:         add $l$ to back_set($l^{\prime\prime}$)
To initialize (Lines 1-3), DMS* begins by creating an initial label $l_{o}$ and invokes the procedure SolveMHPP (as aforementioned) to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$. For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$. Here, $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals. In contrast, $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$. Similarly to A* [30], let OPEN denote a priority queue that stores labels $l$ and prioritizes them based on their $f$-values from the minimum to the maximum. In Alg. 1, we point out which $f$-value (either $f_{temp}$ or $f_{max}$) is used when a label is added to OPEN. Finally, $l_{o}$ is added to $\mathcal{F}(v_{o})$ since $l_{o}$ is non-dominated by any other labels at $v_{o}$, and $l_{o}$ is added to OPEN for future search. [^4] The notion of frontier set in this paper differs from the ones in . In , $\mathcal{F}(v)$ contains only expanded labels, which resembles the closed set in conventional A*, while in this paper, $\mathcal{F}(v)v\in\mathcal{V^{W}}$ includes both the labels that are expanded (closed labels) and labels that are generated and not yet expanded (open labels).
In each search iteration (Lines 5-27), DMS* begins by popping a label $l$ from OPEN for processing. DMS* first calls the TargetSeq procedure for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$. TargetSeq determines either to call SolveMHPP to find a joint sequence for $l$ (i.e., $\gamma(l)$), or to re-use the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$ as $\gamma(l)$. We elaborate TargetSeq in Sec. IV-C. After TargetSeq on Line 7, the heuristic $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes the $f_{max}(l)$ and compare it against $f_{temp}(l)$. When computing $f_{max}$ out of $\vec{g}$ and $\vec{h}$ on Line 8, a heuristic inflation factor $w\in[1,\infty)$ is used, which scales each component in $\vec{h}$ by the factor $w$. Heuristic inflation is a common technique for A* [31] and M*-based algorithms [10] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [31]. If $f_{max}(l)>f_{temp}(l)$, then $l$ should not be expanded in the current iteration since there can be labels in OPEN that have smaller $f$-value than $f_{max}(l)$. DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and skip the current search iteration. In a future search iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
Afterwards, DMS* checks if $l$ leads to a solution using CheckSuccess($l$), which verifies if every component in $\vec{a}(l)$ is one and if every component of $v(l)$ is a unique goal vertex while satisfying the assignment constraints. If CheckSuccess($l$) returns true, a solution joint path $\pi^{*}$ is found and can be reconstructed by iterative tracking the parent pointers of labels from $l$ to $l_{o}$ in Reconstruct($l$). DMS* then terminates.
If CheckSuccess($l$) returns false, $l$ is expanded by considering its limited neighbors [10] described as follows. Let $I_{C}(l)\subseteq I$ denote the conflict set of label $l$, which is a subset of agents that are (or will be) in conflict. The limited neighbors of $l$ is a set of successor labels of $l$. For each agent $i$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$. If $i\in I_{C}(l)$, agent $i$ is allowed to visit any adjacent vertex of $v^{i}(l)$ in $G^{W}$. In other words, the successor vertices of $v^{i}(l)$ are the following:
$u^{i}\leftarrow\begin{cases}\phi^{i}_{\gamma(l)}(v^{i}(l))&\mbox{if }i\notin I_{C}(l)\\ u^{i}\;|\;(v^{i}(l),u^{i})\in E^{W}&\mbox{if }i\in I_{C}(l)\end{cases}$ (1)
Let $V_{succ}^{i}$ denote the set of successor vertices of $v^{i}(l)$, which is either of size one or equal to the number of edges incident on $v^{i}(l)$ in $G^{W}$. The successor joint vertices $V_{succ}$ of $v(l)$ is then the combination of $v^{i}(l)$ for all $i\in I$, i.e., $V_{succ}:=V_{succ}^{1}\times V_{succ}^{2}\times\cdots\times V_{succ}^{N}$. For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding label $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$. When creating $l^{\prime}$, the corresponding $\vec{g}(l^{\prime})$ and $\vec{a}(l^{\prime})$ are computed based on $\vec{g}(l)$ $\vec{a}(l)$ and $v^{\prime}$.
After generating the successor labels $L_{succ}$ of $l$, for each label $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$. DMS* then invokes a procedure BackProp (Alg. 3) to back propagate $I_{C}(l^{\prime})$ to its ancestor labels recursively so that the conflict set of these ancestor labels are modified, and labels with modified conflict set are re-added to OPEN and will be re-expanded. DMS* maintains a back_set($l$) for each label $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted. Intuitively, similarly to [10, 8], the conflict set of labels are dynamically enlarged during planning when agents are detected in conflict. The conflict sets of labels determine the sub-graph within the joint graph $\mathcal{G^{W}}$ that can be reached by DMS*, and DMS* always attempts to limit the search within a sub-graph of $\mathcal{G^{W}}$ as small as possible.
Afterwards, if the conflict set $I_{C}(l^{\prime})$ is non-empty, then label $l^{\prime}$ leads to a conflict and is discarded. Otherwise, $l^{\prime}$ is checked for dominance against any existing labels in $\mathcal{F}(v(l^{\prime}))$ using Def. 3. If $l^{\prime}$ is dominated by or is equal to any existing labels in $\mathcal{F}(v(l^{\prime}))$, $l^{\prime}$ is pruned, since any future joint path from $l^{\prime}$ can be cut and paste to a label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach goals. Furthermore, for each label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, the procedure DomBackProp(Alg. 4) is invoked so that the conflict set $I_{C}(l^{\prime})$ is back propagated to $l$, and $l$ is added to the back_set of $l^{\prime\prime}$. By doing so, DMS* is able to keep updating the conflict set of the predecessor labels of $l^{\prime}$ after $l^{\prime}$ is pruned. This ensures that, when needed, the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned.
If label $l^{\prime}$ is not pruned, a procedure SimpleHeu is invoked for $l^{\prime}$ to quickly compute a heuristic that roughly estimates the cost-to-go. One possible way to implement SimpleHeu is to first copy $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero. This heuristic is still an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion. Then, the $f_{temp}(l^{\prime})$ can be computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority. Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the current search iteration ends.
When DMS* terminates, it either finds a conflict-free joint path, or returns failure when OPEN is empty if the given instance is unsolvable.
IV-C Deferred Target Sequencing
DMS* introduces two techniques to defer the target sequencing until absolutely needed. As aforementioned, the first one uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic when a label is generated, and invokes TargetSeq only when that label is popped from OPEN for expansion.
We now focus on the second technique in DMS*. In the previous MS* [8], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, the procedure SolveMHPP needs to be invoked for $l$ to find a joint sequence and policy from $l$, which makes the overall computation burdensome, especially when a lot of new labels are generated due to the agent-agent conflict.
Different from MS*, DMS* seeks to defer the call of SolveMHPP inside TargetSeq. For a label $l^{\prime}$, DMS* attempts to avoid calling SolveMHPP for $l^{\prime}$ by re-using the joint sequence $\gamma(l)$ of its parent label $l$ ($l$ is the parent of $l^{\prime}$). As presented in Alg. 2, on Lines 5-6, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$. Agents $i$ that are not in the conflict set $I_{C}(l)$ (i.e., $i\notin I_{C}(l)$) are still along their individual paths as specified by the previously computed policy $\phi_{\gamma(l)}$. Agents $i$ that are in the conflict set $I_{C}(l)$ (i.e., $i\in I_{C}(l)$) consider all possible actions as described in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$. Therefore, DMS* needs to go through a check for these agents $i\in I_{C}(l)$: DMS* first computes the $f$-vector by summing up $\vec{g}(l^{\prime})$ and $\vec{h}(l^{\prime})$. Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value. Otherwise, there exists an agent $i\in I_{C}(l)$, whose corresponding $f^{i}(l^{\prime})$ is greater than $f_{max}(l)$, and in this case, DMS* cannot avoid calling SolveMHPP since there may exists another joint sequence from $l^{\prime}$, which leads to a solution joint path with better (smaller) objective value.
IV-D Properties of DMS*
IV-D1 Completeness
This section discusses the properties of DMS*. An algorithm is complete for a problem if finds a solution for solvable instances, and reports failure in finite time for unsolvable instances.
Theorem 1.
DMS* is complete for MCPF-max.
Let $\mathcal{A}=\{0,1\}^{M}$ denote the set of all possible binary vectors and let $\mathcal{S}=\mathcal{G^{W}}\times\mathcal{A}$ denote the Cartesian of the joint graph $\mathcal{G^{W}}$ and $\mathcal{A}$. $\mathcal{S}$ is a finite space. DMS* uses dominance pruning and does not expand the same label twice during the search. As a result, for a unsolvable instance, DMS* enumerates all possible joint paths that starts from $(v_{o},\vec{0}^{M})$ to any reachable $(v,a)\in\mathcal{S}$ and terminates in finite time. For a solvable instance, DMS* conducts systematic search in $\mathcal{S}$ and returns a solution in finite time.
IV-D2 Solution Optimality
To discuss the solution optimality of DMS*, we needs the following assumptions.
Assumption 1.
The procedure SolveMHPP returns an optimal joint sequence for the given mHPP instance.
Assumption 2.
Line 7 in Alg. 1 never returned true during the search of DMS*.
Theorem 2.
For a solvable instance, when Assumption 1 and 2 hold, DMS* returns an optimal solution joint path.
The proof of Theorem 2 follows the analysis in [10, 8], and we highlight the main ideas here. The policies computed by DMS* defines a sub-graph $\mathcal{G}^{W}_{sub}$ of the joint graph $\mathcal{G^{W}}$. DMS* first expands labels in this sub-graph $\mathcal{G}^{W}_{sub}$. If no conflict is detected when following the policy, then the resulting joint path is conflict-free and optimal due to Assumption 1: With Assumption 1, ${h}_{max}(l)$ for any label $l$ computed by SolveMHPP is an estimated cost-to-go that is admissible, i.e., ${h}_{max}(l)$ is no larger than the true optimal cost-to-go. This is true because SolveMHPP ignores agent-agent conflicts and the resulting ${h}_{max}(l)$ must be a lower bound on the true cost-to-go. As DMS* selects labels from OPEN in the same way as A* does, admissible heuristics lead to an optimal solution [31].
If conflicts are detected, DMS* updates (i.e., enlarges) $\mathcal{G}^{W}_{sub}$ by enlarging the conflict set and back propagating the conflict set. When Assumption 2 holds, the enlarged $\mathcal{G}^{W}_{sub}$ still ensures that an optimal conflict-free joint path $\pi_{*}$ is contained in the updated $\mathcal{G}^{W}_{sub}$ [10, 8]. DMS* systematically search over $\mathcal{G}^{W}_{sub}$ and finds $\pi_{*}$ at termination.
We now explain if Assumption 2 is violated, why DMS* loses the optimality guarantee and how to fix it. In MCPF-max, all agents are “coupled” in the space of binary vectors $\mathcal{A}$ in a sense that a target visited by one agent does not need to be visited by any other agents. As as result, when one agent changes its target sequence by visiting some target that is previously assigned to another agent, all agents may need to be re-planned in order to ensure solution optimality. It means that, when Line 7 in Alg. 1 returns true for label $l^{\prime}$, procedure SolveMHPP needs to be called for $l^{\prime}$ and SolveMHPP may return a new joint sequence $\gamma^{\prime}$. In this new sequence $\gamma^{\prime}$, it is possible that agents within $I_{C}(l)$ visits targets that are previously assigned to agents that are not in $I_{C}(l)$. However, DMS* does not let agents outside $I_{C}(l)$ to take all possible actions as defined in Equation (1). As a result, an optimal solution joint path may lie outside $\mathcal{G}^{W}_{sub}$ and DMS* thus loses the solution optimality guarantee when Assumption 2 is violated.
To ensure solution optimality when Assumption 2 is violated, same as in [8], one possible way is to back propagate the entire index set $I=\{1,2,\cdots,N\}$ as the conflict set when calling BackProp. This ensures that DMS* considers all possible actions for all agents when conflicts between agents are detected. However, in practice, this is computationally burdensome when $N$ is large, and limits the scalability of the approach. We therefore omit it from Alg. 1 for clarity.
IV-D3 Solution Bounded Sub-Optimality
A joint sequence $\gamma$ is $\epsilon$-bounded sub-optimal ($\epsilon\geq 0$) if $\max_{i\in I}cost(\gamma^{i})\leq(1+\epsilon)\max_{i\in I}cost(\gamma_{*}^{i})$, where $\gamma_{*}$ is an optimal joint sequence. Similarly, a solution joint path $\pi$ is $\epsilon$-bounded sub-optimal if its objective value $g\leq(1+\epsilon)g_{*}$, where $g_{*}$ is the true optimum of the given instance.
Theorem 3.
For a solvable instance, if (i) SolveMHPP returns a $\epsilon$-bounded sub-optimal joint sequence for any given label $l$ and (ii) Assumption 2 holds, then DMS* returns a $\epsilon$-bounded solution joint path.
Bounded sub-optimal joint sequences lead to inflated heuristic values for labels in DMS*. As DMS* select labels based on $f$-values as A* does, inflated heuristic values lead to bounded sub-optimal solutions for the problem [31].
In addition to using bounded sub-optimal joint sequences, DMS* is also able to use the conventional heuristic inflation technique [31] and provide bounded sub-optimal solution.
Theorem 4.
For a solvable instance, if Line 24 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
V Experimental Results
We implement DMS* in Python, and use Google OR-Tool [^5] https://developers.google.com/optimization/routing/routing_tasks as the mHPP solver, which implements the SolveMHPP procedure that is required by DMS*. Limited by our knowledge on Google OR-Tool, we only consider the following type of assignment constraints, where any agent can visit all targets and goals. We use grid maps from the MAPF benchmark data set [32] and set up a 60-second runtime limit for each test instance, where each test instance contains the starts, targets and goals, which are all from the MAPF benchmark data set [32]. All tests run on a MacBook Pro with a Apple M2 Pro CPU and 16GM RAM. It is known from [10] that, for M*-based algorithms, a slightly inflated heuristic can often help improve the scalability of the approach with respect to the number of agents and we set $w=1.1$ in our tests. We set the number of targets $M=20,40,60,80$ in our tests. We compare DMS*, which includes the two proposed technique to defer target sequencing, against MS*, which is the baseline and does not have these two techniques. In other words, MS* here is a naive adaption of the existing MS* [8] algorithm to solve MCPF-max, by using the aforementioned Google OR-Tool to solve min-max mHPP for target sequencing.
Figure 3: The success rates and runtime of both DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map. DMS* achieves higher success rates than MS* while requiring less runtime on average.
V-A Varying Number of Agents
We first fix the map to be Random 32x32, and vary the number of agents $N=5,10,15,20$. We measure the success rates, the average runtime to solution and the average runtime for target sequencing per instance. The averages are taken over all instances, including both succeeded instances and instances where the algorithm times out. As shown in Fig. 3, DMS* achieves higher success rates and lower runtime than the baseline MS* for all $N$s. As $M$ increases from $20$ to $80$, both algorithms require more runtime for target sequencing. As $N$ increases from $5$ to $20$, agents have higher density and are more likely to run into conflict with each other. As a result, both algorithms time out for more and more instances. Additionally, MS* spends almost all of its runtime in target sequencing, which indicates the computational burden caused by the frequent call of mHPP solver. In contrast, for DMS*, when $N=5,10$, DMS* spends most of the runtime in target sequencing, while as $N$ increases to $15,20$, DMS* spends more runtime in path planning.
Figure 4: The success rates and runtime of both DMS* and MS* (baseline) with varying number of targets in maps of different sizes. DMS* achieves higher success rates than MS* while requiring less runtime on average.
V-B Different Maps
We then fixed the number of agents to $N=10$ and test both algorithms in several maps of different sizes, as shown in Fig. 4. DMS* outperforms MS* in terms of success rates due to the alleviated computational burden for target sequencing. Furthermore, larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other. The Room 32x32 map is relatively more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
V-C Experiments with Mobile Robots
We also performed experiments with Husarion ROSbot 2666https://husarion.com/manuals/rosbot/, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization. We used two such robots to perform tests in two different environments, as shown in Fig. 5 (a) and (b) having 2 and 3 targets, respectively. This test verifies that the paths planned by DMS* are executable on real robots. When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [33]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work. In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
Figure 5: Test settings for ROSbot experiments. (a) and (b) show the real-world setup. (c) and (d) show the rviz window corresponding to (a) and (b) respectively. Intermediate targets are highlighted using pink boxes and the goals using green.
We then compare the performance of DMS* (which minimizes the makespan) and MS* (which minimizes the sum of arrival times) on the mobile robots. Both planners were tested on the same map with two intermediate targets as shown in Fig. 6. We observe that DMS* tends to distribute the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its assigned destination. MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
Figure 6: Comparison of DMS* and MS* on mobile robots. (a) shows the test setting. (b) and (c) show the joint path planned by DMS* and MS* respectively. Both planners return different paths since they optimize different objective functions.
VI Conclusion and Future Work
This paper investigates a min-max variant of Multi-Agent Combinatorial Path Finding problem and develops DMS* algorithm to solve this problem. We analyze the properties of DMS* and test DMS* with up to 20 agents and 80 targets. We also conduct robot experiments to showcase the usage of DMS* on differential-drive robots.
For future work, one can investigate how to incorporate the existing conflict resolution techniques for MAPF into the DMS* to further improve its runtime efficiency. We also note from our robot experiments that, the uncertainty and disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to address the uncertainty and disturbance.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. 2120219 and 2120529. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@
 Abstract
-Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their initial to goal locations, while visiting a set of intermediate target locations in the middle of the paths.
+Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their start to goal locations, while visiting a set of intermediate target locations in the middle of the paths.
 MCPF is challenging as it involves both planning collision-free paths for multiple agents and target sequencing, i.e., solving traveling salesman problems to assign targets to and find the visiting order for the agents.
@@
 We demonstrate the use of DMS* on differential-drive robots.
-Index Terms:
-Path Planning for Multiple Mobile Robots or Agents, Multi-Agent Path Finding, Traveling Salesman Problem
 I Introduction
@@
 This paper considers a generalization of MAPF called Multi-Agent Combinatorial Path Finding (MCPF), where the agents need to visit a pre-specified set of intermediate target locations before reaching their goals.
-MAPF and MCPF arise in applications such as surveillance [1] and logistics [2].
-For instance, factories [3, 4] use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing.
+MAPF and MCPF arise in applications such as logistics [1].
+For instance, factories use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing.
 These robots share a cluttered environment and follow collision-free paths.
 In such settings, MAPF problems and their generalizations naturally arise to optimize operations.
-MCPF is challenging as it involves both collision avoidance among the agents as present in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [5, 6] to specify the allocation and visiting orders of targets for all agents.
+MCPF is challenging as it involves both collision avoidance among the agents as in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [2, 3] to specify the assignment and visiting orders of targets for all agents.
 Both the TSP and the MAPF are NP-hard to solve to optimality [4, 2], and so is MCPF.
-A few methods have been developed [8, 9] to address the challenges in MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times.
+A few methods were developed [5, 6] to address MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times.
 Such a formulation may result in an ensemble of paths where some agents arrive early while others arrive late, which leads to long execution times before all agents finish their paths.
@@
 1), denoted as MCPF-max, where the objective is to minimize the maximum arrival time, which is also called the makespan, of all agents.
-To solve MCPF-max, this paper first adapts our prior MS* algorithm [8], which was first designed for MCPF-sum, to address MCPF-max.
+Figure 1: MCPF-max and MCPF-sum.
+MCPF-max seeks a set of collision-free paths while minimizing the maximum arrival time of the agents.
+The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal.
+To solve MCPF-max, we first adapt our prior MS* algorithm [5], which was designed for MCPF-sum, to address MCPF-max.
 Then, we further develop two new techniques to expedite the planning, and we call the resulting new algorithm Deferred MS* (DMS*).
-Figure 1: MCPF-max problem, which seeks a set of collision-free paths while minimizing the maximum arrival time of all agents.
-The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal.
-The orange target appears on the path of the blue agent, but is not claimed to be visited by the blue agent.
-Specifically, the existing MS* follows the general heuristic search framework (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state.
-MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for agents.
-When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs for the agents to arrive at their goals.
+Specifically, the existing MS* is a heuristic search approach (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state.
+MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for the agents.
+When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs to reach the goals.
 Therefore, the cost of target sequences provides an admissible heuristic to guide the state selection and expansion as in A*.
 Furthermore, MS* leverages the idea in M* [7] to first use the target sequences to build a low-dimensional search space, and then grow this search space by coupling agents together for planning only when collision happens.
-By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques under the heuristic search framework and is able to provide completeness and solution quality guarantees.
+By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques using a heuristic search approach, and provides completeness and solution optimality.
 The first technique developed in this paper is applicable to MS* for both MCPF-max and MCPF-sum.
@@
 To remedy this issue, for each generated successor, we first use a fast-to-compute yet roughly estimated cost-to-go as the heuristic value, and defer calling TSP solver for target sequencing until that successor is selected for expansion.
-The second technique developed in this paper is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum.
+The second technique is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum.
 Since the goal here is to minimize the makespan, during the search, agents with non-maximum arrival time naturally have “margins” in a sense that they can take a longer path without worsening the makespan of all agents.
 We take advantage of these margins to let agents re-use their previously computed target sequences and defer the expensive calls of TSP solvers until the margin depletes.
-To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as real robot experiments.
-The simulation shows that the new techniques in DMS* help triple the success rates and reduce the average runtime to solution.
-The real robot experiments demonstrate that the planned path are executable on real robots, and that DMS* is able to optimize makespan while the existing planners that minimize the sum of arrival times can lead to long makespan.
+To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as a simple real robot experiment.
+The simulation shows that: (i) DMS* finds paths that are up to 50% cheaper than an iterative greedy baseline method, and (ii) the new techniques in DMS* help triple the success rates and reduce the average runtime to solution comparing to MS*.
+The robot experiments show that the planned path are executable, and inspire future work.
 II Related Work
-II-A Multi-Agent Path Finding
-To find collision-free paths for multiple agents, a variety of MAPF algorithms were developed, falling on a spectrum from coupled [11] to decoupled [12], trading off completeness and optimality for scalability.
-In the middle of this spectrum lie the dynamically-coupled methods such as M* [10] and CBS [13], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary in order to resolve agent-agent collision.
-These dynamically-coupled methods were improved and extended in many ways [14, 15, 16, 17].
-II-B Traveling Salesman Problems
-In the presence of multiple intermediate targets, the planner needs to determine both the assignment and visiting order of the targets for the agents.
-For a single agent, the well-known Traveling Salesman Problem (TSP), which seeks a shortest tour that visits every vertex in a graph, is one of the most well-known NP-hard problems [5].
+Multi-Agent Path Finding algorithms fall on a spectrum from coupled [8] to decoupled [9], trading off completeness and optimality for scalability.
+In the middle of this spectrum lie the dynamically-coupled methods such as M* [7] and CBS [10], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary to avoid agent-agent collision.
+Traveling Salesman Problems determine both the assignment and visiting order of the targets for the agents, where there are multiple intermediate targets to visit.
+For a single agent, the Traveling Salesman Problem (TSP) seeks a shortest tour that visits every vertex in a graph, and is a well-known NP-hard problem [2].
 Closely related to TSP, the Hamiltonian Path Problem (HPP) requires finding a shortest path that visits each vertex in the graph from a start vertex to a goal vertex.
 The multi-agent version of the TSP and HPP (denoted as mTSP and mHPP, respectively) are more challenging since the vertices in the graph must be allocated to each agent in addition to finding the optimal visiting order of vertices.
-To simplify the presentation, we refer to all these problems simply as TSPs.
-To solve TSPs, a variety of methods have been developed ranging from exact methods (branch and bound, branch and price) [5] to heuristics [18] and approximation algorithms [19], trading off solution optimality for runtime efficiency.
-This work does not develop new TSPs solvers and leverage the existing ones.
-II-C Target Assignment, Sequencing and Path Finding
-Some recent research combines MAPF with target assignment and ordering [20, 21, 22, 23, 24, 25, 9, 8, 26, 27, 28, 29].
+We refer to all these problems simply as TSPs.
+Different methods have been developed [2, 11, 12] to solve TSPs, trading off solution optimality for runtime efficiency.
+This paper does not develop new TSPs solvers and leverage the existing ones.
+Target Assignment, Sequencing and Path Finding were recently combined in different ways [13, 14, 15, 16, 6, 5, 17, 18].
 Most of them either consider target assignment only (without the need for computing visiting orders of targets) [13, 14, 15], or consider the visiting order only given that each agent is pre-allocated a set of targets [16, 17, 18].
-Our prior work [8, 9] seeks to handle the challenge in target assignment and ordering as well as the collision avoidance in MAPF simultaneously.
-These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max, which seeks to minimize the maximum arrival time.
+Our prior work [5, 6] seeks to handle the challenge in target assignment and ordering and the collision avoidance in MAPF simultaneously.
+These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max.
+We do not extend our prior method CBSS [6] for MCPF-sum to MCPF-max, because CBSS requires solving a min-sum K-best sequencing problem, and it is not obvious how to handle the min-max variant of this K-best sequencing problem.
 III Problem
-Let index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents.
+Let the index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents.
 All agents share a workspace that is represented as an undirected graph $G^{W}=(V^{W},E^{W},c^{W})$, where $W$ stands for workspace.
@@
 In this paper, the cost of an edge is equal to its traversal time, and each edge has a unit cost.
-[^1] In [8, 9], the edge cost can be different from its traversal time: the edge cost can be any positive number while the traversal time is always one unit per edge.
-In this paper, the edge cost are the same as the edge traversal time, which is one for each edge.
+[^1] Here, the edge cost is the same as the edge traversal time, which is one unit for each edge.
+When the traversal time of edges are not unitary, it leads to continuous-time MAPF, and we refer the reader to [19, 20].
 Let the superscript $i\in I$ over a variable denote the specific agent to which the variable belongs (e.g.
@@
 Let $v_{o}^{i},v_{d}^{i}\in V^{W}$ denote the initial (or original) vertex and the goal (or destination) vertex of agent $i$ respectively.
-Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}$ denote the set of target vertices.
-For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.
+Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}\backslash\{V_{o}\cup V_{d}\}$ denote the set of target vertices.
+For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.222 An agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited.
+If a target $v$ is in the middle of the path of $i$ and $i$ does not claim $v$ is visited, then $v$ is not considered as visited.
+A visited target $v$ can appear in the path of another agent.
+When we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target.
+The assignment constraints do not forbid any agent $j\notin f_{A}(v)$ to use $v$ in its path, and only forbid agent $j$ to claim visiting $v$.
 Let $\pi^{i}(v^{i}_{1},v^{i}_{\ell})$ denote a path for agent $i$ between vertices $v^{i}_{1}$ and $v^{i}_{\ell}$, which is a list of vertices $(v^{i}_{1},v^{i}_{{2}},\dots,v^{i}_{\ell})$ in $G^{W}$ with $(v^{i}_{k},v^{i}_{k+1})\in E^{W},k=1,2,\cdots,\ell-1$.
@@
 The second case is an edge conflict $(i,j,e,t)$, where two agents $i,j\in I$ go through the same edge $e$ from opposite directions between times $t$ and $t+1$.
-Definition 1 (MCPF-max Problem).
-The Multi-Agent Combinatorial Path Finding with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited [^2] In MCPF-max, the notion that an agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited.
-In other words, if a target $v$ is in the middle of the path of agent $i$ and agent $i$ does not claim $v$ is visited, then $v$ is not considered as visited.
-Additionally, a visited target $v$ can appear in the path of another agent.
-For the rest of the paper, when we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target.
-at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}cost(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
-Remark 1.
-The MCPF-max is similar to MCPF-sum [9, 8].
-The only difference between these two problems is that: MCPF-sum seeks to minimize the sum of individual path costs while MCPF-max minimizes the maximum of individual path costs.
+Definition 1 (MCPF-max Problem)
+The MCPF with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}g(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
 IV Method
-This section begins with a toy example in Fig.
-2 to illustrate the planning process of DMS*, and provides the explanation in the caption.
-This section then introduces the concepts and notations, explains the pseudo-code and elaborates the technical detail.
+This section begins with an example in Fig.
+2 to illustrate the planning process of DMS*.
+Then, Sec.
+IV-A introduces some concepts that will be used during the search process, before the presentation of the search algorithm in Sec.
+IV-B.
 Figure 2: An illustration of DMS* and related concepts.
-(a) shows the workspace graph $G^{W}$ and a toy problem.
-(b) shows the corresponding target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices.
+(a) shows the workspace graph $G^{W}$.
+(b) shows the target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices.
 (c) shows the workflow of DMS* in Alg.
 1.
-(d) shows that DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1).
+(d) shows DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1).
 This joint sequence $\gamma(l_{o})$ can be converted to a joint path $\pi(l_{o})$ (d2,d3), whose corresponding makespan $f_{max}$ is 9.
 (e) The joint path $\pi(l_{o})$ leads to a policy $\phi_{\gamma(l_{o})}$ that maps a joint vertex to the next joint vertex along $\pi(l_{o})$.
-Since agent-agent conflicts are ignored, when following the policy, agents may run into conflict (e2).
-When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the nearby ancestor labels.
-(f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of the agents that are in conflict (f1,f2,f3).
-During this re-expansion, for each of the successor, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP.
-For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means that each agent can move at most one step towards their goals.
-When either of (f2) and (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and see if the popped label should be expanded or re-added to OPEN for future expansion.
-To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this will worsen the overall makespan.
+Since conflicts are ignored in this policy, agents may run into conflict (e2).
+When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the ancestor labels.
+(f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of all agents in conflict (f1,f2,f3).
+During the re-expansion, for each of the successors, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP.
+For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means each agent can move at most one step towards their goals.
+When either (f2) or (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and checks if the popped label should be expanded or re-added to OPEN for future expansion.
+To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this worsens the makespan.
 For the successor shown in (f2), the resulting makespan is 9 (f4), which is no worse than the previous makespan 9, and DMS* will not call mHPP solver to save computational effort.
-For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find an optimal joint sequence from that successor (f3) to the goals.
+For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find a new joint sequence from that successor (f3) to the goals.
 Finally, with the joint sequence, new policy for the successors can be built and the search continues as in (e).
@@
 IV-A1 Joint Graph
-Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of all agents which is the Cartesian product of $N$ copies of $G^{W}$, where each vertex $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices.
-The joint vertex $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ corresponding to the initial vertices of all agents is called the initial joint vertex.
+Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of the agents which is the Cartesian product of $N$ copies of $G^{W}$, where each $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices.
+Let $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ denote the initial joint vertex, which contains the initial vertices of all the agents.
 Let $\pi(u,v),u,v\in\mathcal{V^{W}}$ denote a joint path, which is a tuple of $N$ (individual) paths of the same length, i.e., $\pi(u,v)=(\pi^{1}(u^{1},v^{1}),\cdots,\pi^{N}(u^{N},v^{N}))$.
-For the rest of this paper, we use a “path” to denote an “individual path”.
+A joint path $\pi$ can be viewed as a list of joint vertices $(v_{0},v_{1},\cdots,v_{\ell})$, and the time for the agents to arrive at each joint vertex $v\in\pi$ is specified by the index of $v$ in $\pi$.
+Given two subsequent joint vertices $u,v$ along a joint path $\pi$, let $\textsl{CheckConflict}(u,v)$ denote a procedure that checks for vertex conflicts at $u,v$ and edge conflicts during the transition from $u$ to $v$ among all pairs of agents, and CheckConflict returns a set $I_{C}\subseteq I$ of agents that are in conflict.
+We use a “path” to denote an “individual path”.
 DMS* searches the joint graph $\mathcal{G^{W}}$ for a joint path that solves the MCPF-max.
 IV-A2 Binary Vector
-For any joint vertex $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g.
-$\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited along the joint path, and we need to differentiate between them during the search.
-We therefore introduce the following concepts.
-First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{t_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list.
+For any $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g.
+$\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited, and we need to differentiate between them.
+First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{u_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list.
 [^3] In this paper, for a vector related to targets (e.g.
@@
 $i$ in $g^{i},v^{i}$) to indicate the element in the vector or joint vertex corresponding to an agent.
-Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if the corresponding target $t_{m}$ is visited by some agent, and $a_{m}=0$ otherwise.
+Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$ by a joint path during the search, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if $u_{m}$ is visited, and $a_{m}=0$ otherwise.
 IV-A3 Label
 Let $l=(v,\vec{a},\vec{g})$ denote a label, where $v$ is a joint vertex, $\vec{a}$ is a binary vector of length $M$, and $\vec{g}$ is a cost vector of length $N$.
-Here, each component $g^{i},i\in I$ corresponds to the path cost of agent $i$.
-Intuitively, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path costs as specified by $\vec{g}$.
-To simplify the presentation, given a label $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$.
-Furthermore, let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$.
+Here, each component $g^{i},i\in I$ is the path cost of agent $i$.
+During the search, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path cost $\vec{g}$.
+Given $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$.
+Let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$.
 To solve the MCPF-max in Def.
-1, the algorithm needs to find a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reach the goals, and $g_{max}(l)$ is the objective value to be minimized.
+1, the planner searches for a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reaches the goals, and $g_{max}(l)$ is the objective value to be minimized.
 IV-A4 Label Comparison
-To compare two labels at the same joint vertex (i.e., to compare the joint paths represented by these two labels), we need to compare both their binary vectors as well as the path costs.
-We introduce the following dominance relationship to compare two binary vectors.
-Definition 2 (Binary Dominance).
+To compare two labels at the same joint vertex, we compare both $\vec{a}$ and $\vec{g}$.
+Definition 2 (Binary Dominance)
 For any two binary vectors $\vec{a}$ and $\vec{b}$, $\vec{a}$ dominates $\vec{b}$ ($\vec{a}\succeq_{b}\vec{b}$), if both the following conditions hold: (i) $\forall m\in\{1,2,\cdots,M\}$, ${a}_{m}\geq{b}_{m}$; (ii) $\exists m\in\{1,2,\cdots,M\}$, ${a}_{m}>{b}_{m}$.
-Intuitively, $\vec{a}$ dominates $\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$.
-In addition, two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
-To compare two labels with the same joint vertex, we define the following dominance rule between them.
-Definition 3 (Label Dominance).
-For any two labels ${l}_{1}$ and ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either of the following two conditions holds: (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$.
-Intuitively, if $l_{1}$ dominates $l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$.
-If $l_{1}$ does not dominate $l_{2}$, $l_{1}$ is then non-dominated by $l_{2}$.
+Intuitively, $\vec{a}\succeq_{b}\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$.
+Two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
+Definition 3 (Label Dominance)
+For two labels ${l}_{1}$, ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; or (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$ holds.
+Intuitively, if $l_{1}\succeq_{l}l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$.
+If $l_{1}$ does not dominate $l_{2}$, $l_{2}$ is then non-dominated by $l_{1}$.
 Any two labels are non-dominated (with respect to each other) if each of them is non-dominated by the other.
-Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if they have the same vertex, binary vector and $g_{max}$ (i.e., $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$).
-Note that there is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Def.
+Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$.
+There is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Sec.
 III seeks to minimize the maximum path cost over the agents.
@@
 IV-A5 Target Sequencing
-Let $\gamma^{i}=\{v^{i}_{o},t_{1},t_{2},\cdots,t_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where $v^{i}_{o}\in V_{o}$ is the initial vertex of agent $i\in I$, each $t_{j}\in V_{t},j=1,2,\cdots,k$ is a target vertex and $v^{i}_{d}\in V_{d}$ is a goal vertex.
-A target sequence $\gamma^{i}$ specifies both the assignment and visiting order of targets for agent $i$.
+Let $\gamma^{i}=\{v^{i}_{o},u_{1},u_{2},\cdots,u_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where each $u_{j},j=1,2,\cdots,k$ is a target vertex (i.e., $u_{j}\in V_{t}$).
 Let $\gamma=\{\gamma^{1},\gamma^{2},\cdots,\gamma^{N}\}$ denote a joint sequence, which specify the assignment and visiting order of all targets for all agents.
-Given a label $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited.
-We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, which is a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$.
-In other words, (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$, the vertex of agent $i$ in $v(l)$, visits a set of unvisited targets $\{t_{m}\}$, where the $m$-th component of $\vec{a}(l)$ is zero, and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ collectively visits all unvisited targets as specified in $\vec{a}(l)$.
+Given $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited.
+We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$: (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$ ($v^{i}(l)$ is not necessarily an origin vertex), visits a set of unvisited targets $\{u_{m}\}$ (where the $m$-th component of $\vec{a}(l)$ is zero), and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ together visits all unvisited targets in $\vec{a}(l)$.
+Intuitively, $\gamma(l)$ is a joint sequence that is meant to complete the joint path represented by $l$.
 IV-A6 Target Graph
-Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in the workspace graph $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$.
-The cost of a target sequence $cost(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$.
-Given a label $l$, to find a joint sequence $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved.
-Let SolveMHPP($l$) denote a procedure that solves the mHPP.
-SolveMHPP($l$) first formulates a target graph $G^{T}=(V^{T},E^{T},c^{T})$, which is computed based on the workspace graph $G^{W}$ as follows.
-The vertex set $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{t_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$.
+Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$.
+The cost of a target sequence $c(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$.
+Given $l$, to find $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved as follows.
+First, a target graph $G^{T}=(V^{T},E^{T},c^{T})$ is created based on $G^{W}$.
+The $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{u_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$.
 $G^{T}$ is fully connected and $E^{T}=V^{T}\times V^{T}$.
-The edge cost in $G^{T}$ of any pair of vertices $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the workspace graph $G^{W}$.
+The edge cost in $G^{T}$ of any pair of $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the $G^{W}$.
 An example $G^{T}$ is shown in Fig.
 2(b).
-The aforementioned target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$ and a joint sequence $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as specified by $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints.
+A target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$, and a $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as in $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints.
 The procedure SolveMHPP$(l)$ can be implemented by various existing algorithms for mHPP.
 IV-A7 Heuristic and Policy
-Given a joint sequence $\gamma(l)$, to simplify the notations, let $h^{i}(l):=cost(\gamma^{i}(l))$ be a heuristic value.
-Since agent-agent conflicts are ignored along the target sequences, the vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$.
-When SolveMHPP($l$) solves the mHPP to optimality, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which can be used as an admissible heuristic for the search.
-For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from the initial vertices to goals for all agents by further extending the joint path represented by $l$.
-Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to label $l$.
-Given a label $l$ and its joint sequence $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ can be built out of $\gamma(l)$, which maps one label to another label as follows.
-First, a joint path $\pi$ can be built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l)$ for any agent $i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$.
-Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from one joint vertex $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$.
-A set of corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ can be built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by considering if $v_{k}$ visits any new targets.
-A set of corresponding cost vectors $\vec{g}_{k}$ can also be computed in a similar way as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$.
-As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l_{k}=(v_{k},\vec{a}_{k},\vec{g}_{k})$ to the next label $l_{k+1}=(v_{k+1},\vec{a}_{k+1},\vec{g}_{k+1})$ along the target sequence $\phi_{\gamma(l)}$.
-Given a vertex $v^{i}$ of agent $i$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$.
+Given $\gamma(l)$, let $h^{i}(l):=c(\gamma^{i}(l))$ be a heuristic value.
+The vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$.
+When SolveMHPP($l$) solves the mHPP to optimality, since conflicts are ignored along the target sequences, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which is an admissible heuristic for the search.
+For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from $v_{o}$ to goals for all agents by further extending the joint path represented by $l$.
+Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to $l$.
+Given $l$ and $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ is built out of $\gamma(l)$, mapping one label to another as follows.
+First, a joint path $\pi$ is built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l),i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$.
+Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$.
+The corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ are built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by checking if $v_{k}$ visits any new targets.
+The corresponding cost vectors $\vec{g}_{k}$ are computed similarly as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$.
+As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l=(v,\vec{a},\vec{g})$ to the next label $l^{\prime}=(v^{\prime},\vec{a}^{\prime},\vec{g}^{\prime})$ along the target sequence $\gamma(l)$.
+For a vertex $v^{i}$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$.
 An example of $\phi^{i}_{\gamma(l)}$ is shown in Fig.
 2(d).
-A label $l$ is called on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$ for some $l^{\prime}$.
-Otherwise, $l$ is called off-policy, which means its next label is unknown yet and a mHPP needs to be solved from $l$ to find the policy $\phi_{\gamma(l)}$.
-We denote $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) as the process of computing the joint sequence, heuristic values and joint policy.
+A label $l$ is on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$.
+Otherwise, $l$ is off-policy, i.e., the next label is unknown and a mHPP needs to be solved for $l$ to find the policy $\phi_{\gamma(l)}$.
+Let $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) denote the process of computing the joint sequence, heuristic values and joint policy.
 IV-B DMS* Algorithm
@@
 12:         continue
-[…]
+13:    if $\textsl{CheckSuccess}(l)$ then
 14:    return Reconstruct($l$)
@@
 16:    for all $l^{\prime}\in L_{succ}$ do
-[…]
+17:         $I_{C}(l^{\prime})\leftarrow\textsl{CheckConflict}(v(l),v(l^{\prime}))$
 18:         BackProp($l$, $I_{C}(l^{\prime})$)
-19:         if $I_{C}\neq\emptyset$ then
-20:             continue
-21:         if IsDominated($l^{\prime}$then
-22:             DomBackProp($l,l^{\prime}$)
-23:             continue
-[…]
-25:         $parent(l^{\prime})\leftarrow l$
-26:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
-27:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
-28:return Failure (no solution)
+19:         if $I_{C}\neq\emptyset$ continue then
+20:         if IsDominated($l^{\prime}$then
+21:             DomBackProp($l,l^{\prime}$)
+22:             continue
+23:         $f_{temp}(l^{\prime})\leftarrow\max_{i\in I}\{\vec{g}(l^{\prime})+w\cdot\textsl{SimpleHeu}(l^{\prime})\}$
+24:         $parent(l^{\prime})\leftarrow l$
+25:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
+26:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
+27:return Failure (no solution)
 Algorithm 2 Pseudocode for TargetSeq($l$, $l^{\prime}$, $I_{C}(l)$)
-1:INPUT: two labels $l,l^{\prime}$ where $l$ is the parent of $l^{\prime}$.
-2:if $l^{\prime}$ is on-policy then
+1:if $l^{\prime}$ is on-policy then
+2:    $\gamma(l^{\prime})\leftarrow\gamma(l)$, return
 3:$\gamma(l^{\prime})\leftarrow\gamma(l)$
-4:    return
-5:$\gamma(l^{\prime})\leftarrow\gamma(l)$
-6:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
-7:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
-[…]
-9:return
+4:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
+5:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
+6:    $\gamma(l^{\prime}),\vec{h}(l^{\prime}),\phi_{\gamma(l^{\prime})}\leftarrow\textsl{SolveMHPP}(l^{\prime})$
+7:return
 Algorithm 3 Pseudocode for BackProp($l,I_{C}(l^{\prime})$)
@@
 2:    $I_{C}(l)\leftarrow I_{C}(l^{\prime})\bigcup I_{C}(l)$
-3:    if $l\notin$ OPEN then
-4:         add $l$ to OPEN
-5:    for all $l^{\prime\prime}\in$ back_set($l$do
-6:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
+3:    if $l\notin$ OPEN then add $l$ to OPEN
+4:    for all $l^{\prime\prime}\in$ back_set($l$do
+5:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
 Algorithm 4 Pseudocode for DomBackProp($l,l^{\prime}$)
@@
 4:         add $l$ to back_set($l^{\prime\prime}$)
-To initialize (Lines 1-3), DMS* begins by creating an initial label $l_{o}$ and invokes the procedure SolveMHPP (as aforementioned) to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$.
-For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$.
-Here, $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals.
-In contrast, $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$.
-Similarly to A* [30], let OPEN denote a priority queue that stores labels $l$ and prioritizes them based on their $f$-values from the minimum to the maximum.
+To initialize (Lines 1-3), DMS* creates an initial label $l_{o}$ and calls SolveMHPP to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$.
+For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$, where $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals; and $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$.
+Similarly to A* [21], let OPEN denote a priority queue storing labels and prioritizing labels based on their $f$-values from the minimum to the maximum.
 In Alg.
@@
 Finally, $l_{o}$ is added to $\mathcal{F}(v_{o})$ since $l_{o}$ is non-dominated by any other labels at $v_{o}$, and $l_{o}$ is added to OPEN for future search.
-[^4] The notion of frontier set in this paper differs from the ones in .
-In , $\mathcal{F}(v)$ contains only expanded labels, which resembles the closed set in conventional A*, while in this paper, $\mathcal{F}(v)v\in\mathcal{V^{W}}$ includes both the labels that are expanded (closed labels) and labels that are generated and not yet expanded (open labels).
-In each search iteration (Lines 5-27), DMS* begins by popping a label $l$ from OPEN for processing.
-DMS* first calls the TargetSeq procedure for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$.
-TargetSeq determines either to call SolveMHPP to find a joint sequence for $l$ (i.e., $\gamma(l)$), or to re-use the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$ as $\gamma(l)$.
+In each iteration (Lines 5-26), DMS* pops a label $l$ from OPEN.
+DMS* calls TargetSeq for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$.
+TargetSeq either calls SolveMHPP to find a joint sequence $\gamma(l)$, or re-uses the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$.
 We elaborate TargetSeq in Sec.
 IV-C.
-After TargetSeq on Line 7, the heuristic $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes the $f_{max}(l)$ and compare it against $f_{temp}(l)$.
+After TargetSeq, $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes $f_{max}(l)$ and compare it against $f_{temp}(l)$.
 When computing $f_{max}$ out of $\vec{g}$ and $\vec{h}$ on Line 8, a heuristic inflation factor $w\in[1,\infty)$ is used, which scales each component in $\vec{h}$ by the factor $w$.
-Heuristic inflation is a common technique for A* [31] and M*-based algorithms [10] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [31].
+Heuristic inflation is common for A* [22] and M*-based algorithms [7] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [22].
 If $f_{max}(l)>f_{temp}(l)$, then $l$ should not be expanded in the current iteration since there can be labels in OPEN that have smaller $f$-value than $f_{max}(l)$.
-DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and skip the current search iteration.
-In a future search iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
+DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and ends the iteration.
+In a future iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
 Afterwards, DMS* checks if $l$ leads to a solution using CheckSuccess($l$), which verifies if every component in $\vec{a}(l)$ is one and if every component of $v(l)$ is a unique goal vertex while satisfying the assignment constraints.
@@
 If CheckSuccess($l$) returns false, $l$ is expanded by considering its limited neighbors [7] described as follows.
-Let $I_{C}(l)\subseteq I$ denote the conflict set of label $l$, which is a subset of agents that are (or will be) in conflict.
 The limited neighbors of $l$ is a set of successor labels of $l$.
-For each agent $i$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$.
+For each $i\in I$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$.
 If $i\in I_{C}(l)$, agent $i$ is allowed to visit any adjacent vertex of $v^{i}(l)$ in $G^{W}$.
-In other words, the successor vertices of $v^{i}(l)$ are the following:
+The successor vertices of $v^{i}(l)$ are:
 $u^{i}\leftarrow\begin{cases}\phi^{i}_{\gamma(l)}(v^{i}(l))&\mbox{if }i\notin I_{C}(l)\\ u^{i}\;|\;(v^{i}(l),u^{i})\in E^{W}&\mbox{if }i\in I_{C}(l)\end{cases}$ (1)
@@
 The successor joint vertices $V_{succ}$ of $v(l)$ is then the combination of $v^{i}(l)$ for all $i\in I$, i.e., $V_{succ}:=V_{succ}^{1}\times V_{succ}^{2}\times\cdots\times V_{succ}^{N}$.
-For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding label $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$.
+For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$.
 When creating $l^{\prime}$, the corresponding $\vec{g}(l^{\prime})$ and $\vec{a}(l^{\prime})$ are computed based on $\vec{g}(l)$ $\vec{a}(l)$ and $v^{\prime}$.
-After generating the successor labels $L_{succ}$ of $l$, for each label $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$.
-DMS* then invokes a procedure BackProp (Alg.
+In other words, the element in $\vec{g}(l^{\prime})$ is one unit larger than the corresponding element in $\vec{g}(l)$ (unless the agent has reached the goal and stays there) since every agents takes an action, and the element in $\vec{a}(l^{\prime})$ changes its value from 0 to 1, if $v^{\prime}$ visits any targets that are unvisited as in $a(l)$.
+After generating the successor labels $L_{succ}$ of $l$, for each $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$.
+DMS* then invokes BackProp (Alg.
 3) to back propagate $I_{C}(l^{\prime})$ to its ancestor labels recursively so that the conflict set of these ancestor labels are modified, and labels with modified conflict set are re-added to OPEN and will be re-expanded.
-DMS* maintains a back_set($l$) for each label $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted.
+DMS* maintains a back_set($l$) for each $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted.
 Intuitively, similarly to [7, 5], the conflict set of labels are dynamically enlarged during planning when agents are detected in conflict.
 The conflict sets of labels determine the sub-graph within the joint graph $\mathcal{G^{W}}$ that can be reached by DMS*, and DMS* always attempts to limit the search within a sub-graph of $\mathcal{G^{W}}$ as small as possible.
-Afterwards, if the conflict set $I_{C}(l^{\prime})$ is non-empty, then label $l^{\prime}$ leads to a conflict and is discarded.
-Otherwise, $l^{\prime}$ is checked for dominance against any existing labels in $\mathcal{F}(v(l^{\prime}))$ using Def.
-3.
-If $l^{\prime}$ is dominated by or is equal to any existing labels in $\mathcal{F}(v(l^{\prime}))$, $l^{\prime}$ is pruned, since any future joint path from $l^{\prime}$ can be cut and paste to a label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach goals.
-Furthermore, for each label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, the procedure DomBackProp(Alg.
+Afterwards, if $I_{C}(l^{\prime})\neq\emptyset$, $l^{\prime}$ leads to a conflict and is discarded.
+Otherwise, $l^{\prime}$ is checked for pruning by using dominance (Def.
+3) against any existing labels in $\mathcal{F}(v(l^{\prime}))$.
+If $l^{\prime}$ is pruned, any future joint path from $l^{\prime}$ can be cut and paste to $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach the goals.
+Furthermore, for each $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, DomBackProp(Alg.
 4) is invoked so that the conflict set $I_{C}(l^{\prime})$ is back propagated to $l$, and $l$ is added to the back_set of $l^{\prime\prime}$.
 By doing so, DMS* is able to keep updating the conflict set of the predecessor labels of $l^{\prime}$ after $l^{\prime}$ is pruned.
-This ensures that, when needed, the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned.
-If label $l^{\prime}$ is not pruned, a procedure SimpleHeu is invoked for $l^{\prime}$ to quickly compute a heuristic that roughly estimates the cost-to-go.
-One possible way to implement SimpleHeu is to first copy $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero.
-This heuristic is still an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion.
-Then, the $f_{temp}(l^{\prime})$ can be computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority.
-Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the current search iteration ends.
+This ensures the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned.
+If $l^{\prime}$ is not pruned, SimpleHeu is invoked for $l^{\prime}$ to compute a heuristic, which can be implemented by first copying $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero.
+This heuristic is an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion.
+Then, $f_{temp}(l^{\prime})$ is computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority.
+Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the iteration ends.
 When DMS* terminates, it either finds a conflict-free joint path, or returns failure when OPEN is empty if the given instance is unsolvable.
 IV-C Deferred Target Sequencing
-DMS* introduces two techniques to defer the target sequencing until absolutely needed.
+DMS* introduces two techniques to defer the target sequencing until needed.
 As aforementioned, the first one uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic when a label is generated, and invokes TargetSeq only when that label is popped from OPEN for expansion.
 We now focus on the second technique in DMS*.
-In the previous MS* [8], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, the procedure SolveMHPP needs to be invoked for $l$ to find a joint sequence and policy from $l$, which makes the overall computation burdensome, especially when a lot of new labels are generated due to the agent-agent conflict.
+In the previous MS* [5], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, SolveMHPP is invoked for $l$ to find a joint sequence and policy from $l$, which burdens the overall computation, especially when a lot of new labels are generated due to the agent-agent conflict.
 Different from MS*, DMS* seeks to defer the call of SolveMHPP inside TargetSeq.
@@
 As presented in Alg.
-2, on Lines 5-6, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$.
-Agents $i$ that are not in the conflict set $I_{C}(l)$ (i.e., $i\notin I_{C}(l)$) are still along their individual paths as specified by the previously computed policy $\phi_{\gamma(l)}$.
-Agents $i$ that are in the conflict set $I_{C}(l)$ (i.e., $i\in I_{C}(l)$) consider all possible actions as described in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$.
+2, on Lines 3-4, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$.
+Agents $i\notin I_{C}(l)$ are still along their individual paths as in the previously computed policy $\phi_{\gamma(l)}$.
+Agents $i\in I_{C}(l)$ consider all actions as in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$.
 Therefore, DMS* needs to go through a check for these agents $i\in I_{C}(l)$: DMS* first computes the $f$-vector by summing up $\vec{g}(l^{\prime})$ and $\vec{h}(l^{\prime})$.
-Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value.
+Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value, the makespan.
 Otherwise, there exists an agent $i\in I_{C}(l)$, whose corresponding $f^{i}(l^{\prime})$ is greater than $f_{max}(l)$, and in this case, DMS* cannot avoid calling SolveMHPP since there may exists another joint sequence from $l^{\prime}$, which leads to a solution joint path with better (smaller) objective value.
-IV-D Properties of DMS*
-IV-D1 Completeness
-This section discusses the properties of DMS*. An algorithm is complete for a problem if finds a solution for solvable instances, and reports failure in finite time for unsolvable instances.
-Theorem 1.
-DMS* is complete for MCPF-max.
+V Analysis
+V-A Completeness
+This section discusses the properties of DMS*. An algorithm is complete for a problem if it finds a solution for any solvable instance, and reports failure in finite time for unsolvable instances.
+Theorem 1
+If SolveMHPP is complete, then DMS* is complete for MCPF-max.
+Proof:
 Let $\mathcal{A}=\{0,1\}^{M}$ denote the set of all possible binary vectors and let $\mathcal{S}=\mathcal{G^{W}}\times\mathcal{A}$ denote the Cartesian of the joint graph $\mathcal{G^{W}}$ and $\mathcal{A}$.
 $\mathcal{S}$ is a finite space.
-DMS* uses dominance pruning and does not expand the same label twice during the search.
+DMS* uses dominance pruning and does not expand the same label twice during the search (the same label will be pruned by dominance).
+The policies computed by DMS* defines a sub-graph $\mathcal{G}^{W}_{sub}$ of the joint graph $\mathcal{G^{W}}$.
+DMS* first expands labels in this sub-graph $\mathcal{G}^{W}_{sub}$.
+If no conflict is detected when following the policy, then the resulting joint path is conflict-free and solves the problem.
+If there is a conflict, DMS* may re-expand a label with a larger conflict set.
+The conflict set of a label can be enlarged for at most $(N-1)$ times.
+Therefore, there is only a finite number of re-expansion of labels.
 As a result, for a unsolvable instance, DMS* enumerates all possible joint paths that starts from $(v_{o},\vec{0}^{M})$ to any reachable $(v,a)\in\mathcal{S}$ and terminates in finite time.
 For a solvable instance, DMS* conducts systematic search in $\mathcal{S}$ and returns a solution in finite time.
-IV-D2 Solution Optimality
+∎
+The above analysis requires that the procedure SolveMHPP is complete for MHPP.
+Otherwise, if SolveMHPP never terminates for certain MHPP instances, DMS* will not terminate in finite time when calling SolveMHPP with those MHPP instances.
+V-B Solution Optimality
 To discuss the solution optimality of DMS*, we needs the following assumptions.
-Assumption 1.
+Assumption 1
 The procedure SolveMHPP returns an optimal joint sequence for the given mHPP instance.
-Assumption 2.
-Line 7 in Alg. 1 never returned true during the search of DMS*.
-Theorem 2.
+Assumption 2
+Line 5 in Alg. 1 never returned true during the search of DMS*.
+Theorem 2
 For a solvable instance, when Assumption 1 and 2 hold, DMS* returns an optimal solution joint path.
@@
 As as result, when one agent changes its target sequence by visiting some target that is previously assigned to another agent, all agents may need to be re-planned in order to ensure solution optimality.
-It means that, when Line 7 in Alg.
+It means that, when Line 5 in Alg.
 1 returns true for label $l^{\prime}$, procedure SolveMHPP needs to be called for $l^{\prime}$ and SolveMHPP may return a new joint sequence $\gamma^{\prime}$.
@@
 1 for clarity.
-IV-D3 Solution Bounded Sub-Optimality
+V-C Solution Bounded Sub-Optimality
 A joint sequence $\gamma$ is $\epsilon$-bounded sub-optimal ($\epsilon\geq 0$) if $\max_{i\in I}cost(\gamma^{i})\leq(1+\epsilon)\max_{i\in I}cost(\gamma_{*}^{i})$, where $\gamma_{*}$ is an optimal joint sequence.
 Similarly, a solution joint path $\pi$ is $\epsilon$-bounded sub-optimal if its objective value $g\leq(1+\epsilon)g_{*}$, where $g_{*}$ is the true optimum of the given instance.
-Theorem 3.
+Theorem 3
 For a solvable instance, if (i) SolveMHPP returns a $\epsilon$-bounded sub-optimal joint sequence for any given label $l$ and (ii) Assumption 2 holds, then DMS* returns a $\epsilon$-bounded solution joint path.
+Proof:
 Bounded sub-optimal joint sequences lead to inflated heuristic values for labels in DMS*.
 As DMS* select labels based on $f$-values as A* does, inflated heuristic values lead to bounded sub-optimal solutions for the problem [22].
+∎
 In addition to using bounded sub-optimal joint sequences, DMS* is also able to use the conventional heuristic inflation technique [22] and provide bounded sub-optimal solution.
-Theorem 4.
-For a solvable instance, if Line 24 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
-V Experimental Results
-We implement DMS* in Python, and use Google OR-Tool [^5] https://developers.google.com/optimization/routing/routing_tasks as the mHPP solver, which implements the SolveMHPP procedure that is required by DMS*.
+Theorem 4
+For a solvable instance, if Line 23 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
+VI Experimental Results
+We implement DMS* in Python, with Google OR-Tool as the mHPP solver.
 Limited by our knowledge on Google OR-Tool, we only consider the following type of assignment constraints, where any agent can visit all targets and goals.
-We use grid maps from the MAPF benchmark data set [32] and set up a 60-second runtime limit for each test instance, where each test instance contains the starts, targets and goals, which are all from the MAPF benchmark data set [32].
+We set up a 60-second runtime limit for each instance, where each instance contains the starts, targets and goals in a grid map.
 All tests run on a MacBook Pro with a Apple M2 Pro CPU and 16GM RAM.
-It is known from [10] that, for M*-based algorithms, a slightly inflated heuristic can often help improve the scalability of the approach with respect to the number of agents and we set $w=1.1$ in our tests.
-We set the number of targets $M=20,40,60,80$ in our tests.
-We compare DMS*, which includes the two proposed technique to defer target sequencing, against MS*, which is the baseline and does not have these two techniques.
+It is known from [7], M*-based algorithms with inflated heuristics can often scale to more agents, and we set $w=1.1$ in our tests.
+We set the number of targets $M=20,40,60,80$.
+We compare DMS*, which includes the two proposed technique to defer target sequencing, against two baselines.
+The first one is MS*, which does not have these two techniques.
 In other words, MS* here is a naive adaption of the existing MS* [5] algorithm to solve MCPF-max, by using the aforementioned Google OR-Tool to solve min-max mHPP for target sequencing.
-Figure 3: The success rates and runtime of both DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map.
+The second baseline is a iterative greedy approach which assigns one unvisited target to an agent to minimize the makespan of all agents in the current iteration while planning collision free paths.
+Figure 3: The success rates and runtime of DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map. DMS* has higher success rates and less runtime on average than MS*.
+Figure 4: The success rates and runtime of both DMS* and MS* (baseline) with varying number of targets in maps of different sizes.
 DMS* achieves higher success rates than MS* while requiring less runtime on average.
-V-A Varying Number of Agents
-We first fix the map to be Random 32x32, and vary the number of agents $N=5,10,15,20$.
+VI-A Simulation Results
+VI-A1 Varying Number of Agents
+We first fix the map to Random 32x32, and vary the number of agents $N=5,10,15,20$.
 We measure the success rates, the average runtime to solution and the average runtime for target sequencing per instance.
@@
 As shown in Fig.
-3, DMS* achieves higher success rates and lower runtime than the baseline MS* for all $N$s.
+3, DMS* achieves higher success rates and lower runtime than the baseline MS*.
 As $M$ increases from $20$ to $80$, both algorithms require more runtime for target sequencing.
 As $N$ increases from $5$ to $20$, agents have higher density and are more likely to run into conflict with each other.
-As a result, both algorithms time out for more and more instances.
+As a result, both algorithms time out for more instances.
 Additionally, MS* spends almost all of its runtime in target sequencing, which indicates the computational burden caused by the frequent call of mHPP solver.
 In contrast, for DMS*, when $N=5,10$, DMS* spends most of the runtime in target sequencing, while as $N$ increases to $15,20$, DMS* spends more runtime in path planning.
-Figure 4: The success rates and runtime of both DMS* and MS* (baseline) with varying number of targets in maps of different sizes.
-DMS* achieves higher success rates than MS* while requiring less runtime on average.
-V-B Different Maps
-We then fixed the number of agents to $N=10$ and test both algorithms in several maps of different sizes, as shown in Fig.
-4.
-DMS* outperforms MS* in terms of success rates due to the alleviated computational burden for target sequencing.
-Furthermore, larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other.
-The Room 32x32 map is relatively more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
-V-C Experiments with Mobile Robots
-We also performed experiments with Husarion ROSbot 2666https://husarion.com/manuals/rosbot/, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization.
-We used two such robots to perform tests in two different environments, as shown in Fig.
-5 (a) and (b) having 2 and 3 targets, respectively.
-This test verifies that the paths planned by DMS* are executable on real robots.
-When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [33]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work.
-In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
+VI-A2 Different Maps
+We then fixed $N=10$ and test in maps of different sizes (Fig.
+4).
+DMS* outperforms MS* in success rates due to the alleviated computational burden for target sequencing.
+Larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other.
+The Room 32x32 map is more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
+$M$ min. CR median CR max.CR
+$20$ 0.47 0.71 0.94
+$40$ 0.46 0.63 0.86
+$60$ 0.53 0.66 0.76
+$80$ 0.56 0.72 0.87
+TABLE I: Cost ratios (CR) of DMS* over a greedy baseline. Solution costs of DMS* are up to 50% cheaper than the greedy.
+VI-A3 Solution Quality
+We use the random 32x32 map and fixed the number of agents to $N=5$ where DMS* achieves 100% success rate.
+We compare the solution cost ratio of DMS* over the greedy baseline.
+As shown in Table I, the solution of DMS* is up to around 50% cheaper than the solution of this greedy baseline.
+$M$ 20 40 60 80
+MS* 72.2 67.6 75.0 71.8
+DMS* 77.2 84.4 93.4 99.9
+TABLE II: The average number of expansion of MS* and DMS* among the instances where both algorithms succeed.
+VI-A4 Number of Expansions
+Table II shows the average number of label expansion in both MS* and DMS* among the instances in the Random 32x32 map which are successfully solved by both algorithms within the runtime limit.
+Due to the deferred sequencing, DMS* tends to search in a less informed manner and needs more expansion.
+VI-B Experiments with Mobile Robots
 Figure 5: Test settings for ROSbot experiments.
-(a) and (b) show the real-world setup.
+(a) and (b) show the setup.
 (c) and (d) show the rviz window corresponding to (a) and (b) respectively.
-Intermediate targets are highlighted using pink boxes and the goals using green.
-We then compare the performance of DMS* (which minimizes the makespan) and MS* (which minimizes the sum of arrival times) on the mobile robots.
-Both planners were tested on the same map with two intermediate targets as shown in Fig.
-6.
-We observe that DMS* tends to distribute the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its assigned destination.
-MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
+Targets are highlighted using pink boxes and the goals using green.
 Figure 6: Comparison of DMS* and MS* on mobile robots.
@@
 Both planners return different paths since they optimize different objective functions.
-VI Conclusion and Future Work
+We also performed experiments with Husarion ROSbot 2, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization.
+We used two such robots to perform tests in two different environments, as shown in Fig.
+5 (a) and (b) having 2 and 3 targets, respectively.
+This test verifies that the paths planned by DMS* are executable on robots.
+When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [23]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work.
+In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
+We then compare the performance of DMS* and MS* on the robots.
+Both planners were tested on the same map with two intermediate targets as shown in Fig.
+6.
+We observe that DMS* tends to assign the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its goal.
+MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
+VII Conclusion and Future Work
 This paper investigates a min-max variant of Multi-Agent Combinatorial Path Finding problem and develops DMS* algorithm to solve this problem.
-We analyze the properties of DMS* and test DMS* with up to 20 agents and 80 targets.
-We also conduct robot experiments to showcase the usage of DMS* on differential-drive robots.
-For future work, one can investigate how to incorporate the existing conflict resolution techniques for MAPF into the DMS* to further improve its runtime efficiency.
-We also note from our robot experiments that, the uncertainty and disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to address the uncertainty and disturbance.
-Acknowledgments
-This material is based upon work supported by the National Science Foundation under Grant No.
-2120219 and 2120529.
-Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
+We test DMS* with up to 20 agents and 80 targets and conduct simple robot experiments to showcase the usage of DMS*.
+For future work, one can investigate extending DMS* to the case where edges have non-unitary traversal times [19, 20], or simultaneously optimizing both the maximum and the sum of individual arrival times [24].
+We note from our experiments the disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to handle the disturbance.